OpenStudy (anonymous):

how do i work out the x intercept of 5x^2log(e)2x

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf y=5x^2log(e)2x}\) \(\Huge\color{blue}{ \sf 5x^2log(e)2x=0}\) Solve for x.

OpenStudy (anonymous):

yeah thats what i need help with

OpenStudy (anonymous):

i forgot to say that its a function

OpenStudy (dumbcow):

something seems wrong about that notation

OpenStudy (anonymous):

if you mean the 2x yeah it should be in brackets

OpenStudy (solomonzelman):

\(\Huge\color{blueviolet}{ \sf 5x^2log(e)2x=0}\) I would combine 5x^2 and 2x first \(\Huge\color{blueviolet}{ \sf 10x^3log(e)=0}\) \(\Huge\color{blueviolet}{ \sf x^3log(e^{10})=0}\) no, I am doing something wrong. Idk sorry.

OpenStudy (dumbcow):

so its \[5x^2 \ln(2x) = 0\]

OpenStudy (anonymous):

yeah

OpenStudy (dumbcow):

5x^2 = 0 when x = 0 ln(2x) = 0 when 2x = 1 ---> x = 1/2

OpenStudy (anonymous):

thars correct answer